Crew 274 Journalist Report 2-7-2023
Author: Tony DiBernardo, Crew Journalist
Crew 274 woke up this morning at 7:30am to the song “There’s a Great Big Beautiful Tomorrow” from the Carousel of Progress at Walt Disney World. The crew got their surveys done, then prepared for a double EVA day utilizing 6 of the 8 crew members.
Today, the crew conducted two EVAs. The purpose of EVA #8 was to deliver a resupply of 2L of water and 6 gel packs to the supply cache, EEVAC-1. HSO Pender collected performance metrics collection and a status check on the condition of the cache and collected data on the orange radiation experiment led by Commander Sarah Guthrie. EEVAC-1 was found in good condition, however, the solar panels charging the cache had an accumulation of dried mud on them. Additionally, most of the legs supporting the solar panels had shifted, and one of them completely collapsed. According to historical weather data, it appears that winds up to 26 mph occurred in the Hanksville area the night of Sol 7. It is suspected that this weather event contributed to the state we found the solar panels in. We were able to clear the mud accumulation with gentle hand swipes and reposition the solar panel legs back to their original configuration. It is noteworthy that the stakes and ropes securing the solar panels all held up over this period, and we believe the stakes installed at the base of the solar panels played a significant role in preventing more drastic impacts.
Additionally, the crew completed the first in a series of extravehicular activities (EVAs) focused on evaluating contingency methodologies and assistive devices, or rather, how we might rescue an incapacitated astronaut on the Moon or Mars while wearing spacesuits. This study uses a Kinetic Utilization Research Tool (“K.U.R.T.”), a weighted grappling dummy; a medical evacuation sled, and an engineered vest with multiple hoist points.
The concept is based on current operation considerations for Artemis missions to the Moon requiring proper suit interface and methods for contingency EVAs. The principal investigator, Commander Guthrie, has performed this study at other analogs, but through that experience, has modified and improved the concept through lessons learned. The unique topography of the Utah desert here at MDRS provides this study with the most analogous environment to perform these tests, which could not be achieved at other analog locations.The crew looks forward to more opportunities in the coming week to rescue “KURT” on the Martian terrain.
Hab Specialist O’Hara participated in EVA 9 this afternoon where he simulated methods for recovering an incapacitated crew member in the field. After the EVA, Bill was able to complete a review of the lower level of the Habitat module. He is on track to complete his study before the conclusion of the mission.
Today, crew astronomers Pena and Loy observed detailed heliophysics phenomena. Multiple observations, videos, and photos were taken of solar dark spots, solar chromosphere granules, and solar prominences. Over 13 large solar prominences were recorded today, the largest dataset of prominences since experimentations began. These observations will be very useful for the astronomer’s analysis of solar cycle 25, how space weather implicates orbital assets, space operations planning to safeguard those assets, and the importance of having heliophysic astronomers on site at Martian Habitats.
Cloudy weather conditions and robotic observatory mechanical failures blocked the astronomy crew from from observing HADS Variable Star V0799 AUR again. Crew Astronomers instead continued their analysis of their current image inventory, stacking and analyzing 9 separate 45-second images in the visual filter. So far, HADS Variable Star V0799 AUR is showing signs of dimming, which provides good data to the astronomers’ variable star report.
No significant progress was seen from any seeds or microgreens today. GreenHab Officer Hines watered the plants in the GreenHab and continued his daily responsibilities.